You are evaluating an investment that will pay ​$80 in 1​ year, and it will continue to make payments at annual intervals​ thereafter, but the payments will grow by 5% forever.

a. What is the present value of the first ​$80 payment if the discount rate is 12%?

b. How much cash will this investment pay 100 years from​ now? What is the present value of the 100th​ payment? Again use a 12​% discount rate.

c. What is the present value of the entire growing stream of perpetual cash​ flows?

d. Explain why the answers to parts a and b help to explain why an infinite stream of growing cash flows has a finite present value.
Answer #1

Present value of a cash flow received after n years at discount rate r is given as:

small PV = rac{FV}{(1+r)^n}

a)

Payment after 1 year is $80 at discount rate 12%, therefore

FV 80 80 PV (1 + r) _ (1+0.12T-1.12 = 71.43

b)

Value of investment after 100 years = 80 + 80*(1+0.05) + 80*(1+0.05)2 + .......... + 80*(1+0.05)99

   = 80 + 80*1.05 + 80*(1.05)2 + .............+ 80*(1.05)99

S0(1.05)100-2082.01 1.051

Sum of a finite GP with n terms and first term a, common ration r = a(rn - 1)/(r-1) when r > 1

Present value of the 100th payment:

80 ((1.05)980125.24 (1+0.12100= 83522.27 = 0.12 PVー

c) Future value of the growing stream = 80 + 80*(1+0.05) + 80*(1+0.05)2 + .......... + 80*(1+0.05)99

Present discounted value

80 80(1 0.05) 80(10.05)2 1 +0.12)(1+0.121 80(1 0.05)99 10.12)100

This is a finite GP of n terms with first term = 80/1.12 = 71.43 and common ratio r = 1.05/1.12 = 0.94

Sum of GP = Present discounted value of growing stream

71.43(1 -(0.94)91187.9 1-0.94

d)

From part (a) and (b), we know that present value of the payment after nth period keeps on decreasing as n increases. So even if payment stream is for infinite years, its present value keeps on decreasing and there is very less contribution from later terms. Hence this stream has a finite present value.